A 200 g whistle firecracker is set and fired horizontally in a straight line on top a smooth platform.

Its speed is 1.654 m/s in 0.35 seconds.

a) What is the momentum ?

b) The force in that time.


1
Expert's answer
2022-06-13T08:23:57-0400

Given:

"v_i=0\:\rm m/s"

"v_f=1.654\:\rm m/s"

"m=0.2\:\rm kg"

"\Delta t=0.35\:\rm s"


(a)

"p_i=mv_i=0.2*0=0\:\rm kg\cdot m/s"

"p_f=mv_f=0.2*1.654=0.33\:\rm kg\cdot m/s"

(b)

"F=\frac{\Delta p}{\Delta t}"

"F=\frac{0.33-0}{0.35}=0.945\:\rm N"